At a university, the correlation coefficient (r) between SAT scores and college GPA is 0.8. Samantha finds the Z score for her SAT score, and wants to use it to predict her Z score for her college GPA. Her SAT Z score is 2. What would be her Z score for her GPA?
Answer #1

here as we know that zy =r*zx ; where zx and zy are z score of independent and dependent variable

therefore Z score for her GPA =0.8*2 =1.6